OpenStudy (anonymous):

Paul went to a solar panel store and paid $923.40 per solar panel for two solar panels and a flat rate fee for installation for a total of $1944.80. Brad went to another solar panel store and paid $940.50 per solar panel for two solar panels plus a per solar panel installation fee for a total of $1983.60. What is the difference in the amounts they paid in installation charges?

OpenStudy (anonymous):

38.8

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

Start with Paul: He paid $923.40 per panel for two panels for a total of $923.40 * 2 = $1846.8 Since he paid $1944.80 total the install fee must be $1944.80-$1846.6 = $98.20 Doing the same thing for Brad we get I=1983.60-(2*940.50)=102.60 Subtracting the two gives us the difference: $102.6-$98.20=$4.40 The difference in the install fess was $4.40

OpenStudy (anonymous):

k yea that would be right dont LISTEN to my answer

OpenStudy (anonymous):

put his down

OpenStudy (anonymous):

OK THANK YOU